what are the spectator ions of lithium sulphate and strontium hydroxide


1
Expert's answer
2021-06-22T04:42:44-0400

Spectator ions are ions that are existing throughout a reaction but are unaffected by it, therefore they are in the same state on both sides of the equation. Without the spectator ions, the equation becomes: H+ ( a q ) + O H- ( a q ) =H2O ( l ). Every neutralisation reaction undergoes this transformation.

Equation

Sr(OH)2 + Li2SO4 → SrSO4 + 2LiOH

so spectator ions for:

lithium sulphate (Li2SO4) are = Li+ and SO4-

and strontium hydroxide (Sr(OH)2) are = Sr2+ and OH-